Calculate the weight  of water in a 1.5 liters 95% nitric acid (HNO3) that has a specific gravity  of 1.82.  Express the concentration of solution in terms of Molarity, Normality and molality.        



1
Expert's answer
2021-02-23T04:25:40-0500

1) density = m/V

Mass of HNO3 = density x V = 1.82 x 1.5 = 2.73 kg

% = (100 x m(solvent))/ (m(solvent) + m(solute))

m(HNO3) = 95/100 x 2.73 = 2.59 kg

m(H2O) = 2.73 - 2.59 = 0.14 kg


2) CM = n/V

M (HNO3) = 63.01 g/mol

n (HNO3) = m/M = 2590/63.01 = 41.10 mol

СM = 41.10/1.5 = 27.4 mol/L


3) Cn = N/V

N = n/equivalence factor

n (HNO3) = 41.10 mol

Equivalence factor (HNO3) = 1

N (HNO3) = 41.10/1 = 41.10

Cn (HNO3) = 41.10/1.5 = 27.4 eq/L


4) Cm = n(solute)/m(solvent)

Cm (HNO3) = 41.10/0.14 = 292.93 mol/kg